Exercise 4.26

Answers

(a)
Let Z represent the number of flips until both Nick and Penny flip Heads. Then is Z ∼FS(p1p2), since Nick’s and Penny’s flips are independent. E(Z) = 1 p1p2
(b)
The logic is analogous to part a, but success probability is p1 + p2 − p1p2.
(c)
P(X1 = X2) = ∑ k=1∞(((1 − p)2)k−1p2) = p 2 − p

(d)
By symmetry, P(X1 < X2) = 1 − ( p 2−p) 2 = 1 − p 2 − p

User profile picture
2021-12-05 00:00
Comments